One angle of a triangle measures 10 degrees more than the smallest angle, and the third angle measures 10 degrees less than twice the smallest angle. Find the measures of the angles.
Answer by cleomenius(959)

Angle A=x
Angle B=x + 10
Angle C=2x - 10
===================================
x + x + 10 + 2x - 10 = 180 degrees
4x = 180 degrees
x = 45 degrees
Angle A=45 degrees
Angle B=55 degrees
Angle C=80 degrees
=================================
45 + 55 + 80 = 180 degrees.
This does check.
Cleomenius.